OpenStudy (anonymous):

Simplify the expression: 3/x + 2/x2

OpenStudy (solomonzelman):

\[\frac{3}{x}+\frac{2}{x^2}~~~~~~~~~~~~ri ght?\]

OpenStudy (anonymous):

I dont know thats why im asking lol

OpenStudy (anonymous):

Is the 2 on the second fraction to the power?

OpenStudy (solomonzelman):

\[\frac{3}{x}+\frac{2}{x^2}\]\[\frac{3 \times x}{x \times x}+\frac{2}{x^2}\]\[\frac{3x}{x^2}+\frac{2}{x^2}\]\[\frac{3x+2}{x^2}\]

OpenStudy (anonymous):

try wolfram

OpenStudy (solomonzelman):

@primeavlqueen22 , see how I did it?

OpenStudy (anonymous):

um kinda lol